# VICON: Vision In-Context Operator Networks for Multi-Physics Fluid Dynamics Prediction

### v4

@@ D.2 Analysis of Checkerboard Artifacts in PDEBench-Comp-HighVis
Training-inference normalization mismatch. During training, context pairs are sampled uniformly across all frames. During inference, context pairs are sampled from initial frames only. As shown in table˜10, the standard deviation at these two different time ranges are vastly different which results in normalization mismatch and exacerbates prediction errors in the already-challenging problem. Future work could explore using global statistics from multiple trajectories instead of single-trajectory instance normalization during deployment. Other potential remedies include using overlapping patches or moving patch positions during rollout, though these represent engineering improvements orthogonal to our core contributions.
D.3 More Results and Visualizations
@@
Step 1
[1e-2]
PDEArena-Incomp 11.01 11.01 5.97 6.17
@@
Step 5
[1e-2]
PDEArena-Incomp 22.68 20.62 11.70 14.81
@@
Step 10
[1e-2]
PDEArena-Incomp 35.67 30.53 20.74 27.89
@@
Last step
[1e-2]
PDEArena-Incomp 76.77 68.03 65.27 93.52
@@
All average
[1e-2]
PDEArena-Incomp 56.27 48.50 41.20 55.95
@@
Step 1
[1e-2]
PDEArena-Incomp 5.63 5.63 3.02 3.12
@@
Step 5
[1e-2]
PDEArena-Incomp 10.38 9.46 5.37 6.79
@@
Step 10
[1e-2]
PDEArena-Incomp 14.65 12.55 8.55 11.49
@@
Last Step
[1e-2]
PDEArena-Incomp 16.50 14.56 14.03 19.47
@@
All average
[1e-2]
PDEArena-Incomp 16.26 14.31 11.86 15.77
@@
Tables 11 and 12 summarize the relative and absolute $L^{2}$ rollout errors across different timesteps for all evaluated models on the three datasets.
@@
fig.˜5 illustrates VICON’s performance with varying timestep strides ($s_{\text{max}}=1,3,5,7$) across all three datasets, demonstrating the model’s ability to adapt to different temporal resolutions without retraining.
@@
fig.˜6 displays the visualization of TKE fields for ground truth and model predictions, highlighting VICON’s superior ability to preserve critical turbulent flow structures compared to baseline approaches.
@@
Table 13: Summary of Rollout Relative $L^{2}$ Error Metrics (single step, scale by std) for 2 Training Strategies. The best results are highlighted in bold. For seperate training (a model on each dataset), each run’s batch size is controlled to be the same as joint training. To ensure fair comparison, we maintain the same batch sizes as joint training for each seperate training runs while adjusting their training steps to be slightly more than one-third of the joint training duration, ensuring comparable total computational costs.
Rollout Relative $L^{2}$ Error [1e-2] Case Joint Seperate
Step 1
PDEArena-Incomp 11.10 9.66
@@
Step 5
PDEArena-Incomp 23.00 24.09
@@
Step 10
PDEArena-Incomp 36.18 41.66
@@
Last Step
PDEArena-Incomp 77.81 123.0
@@
All average
PDEArena-Incomp 56.26 76.35
@@
fig.˜7 and table˜13 quantify the performance differences between jointly trained and separately trained models, demonstrating the significant advantages of multi-physics training across all datasets.
@@
table˜14 summarizes the computational resources required by each method, showing VICON’s advantages in terms of training cost, inference speed, and model parameter count compared to both DPOT and MPP baselines.
@@
fig.˜8 show the rollout results with imperfect temporal measurements.
Figure 9: Comparing outputs from different models. The target is the first 6 output steps from PDEBench-Comp-HighVis dataset (x-velocity channel). For each model(each row), we display the difference between target and model output. Errors (rescale by std) for the full trajectory are listed after the model names.
Figure 10: Comparing outputs from different models. The target is the first 6 output steps from PDEBench-Comp-LowVis dataset (pressure channel). For each model(each row), we display the difference between target and model output. Errors (rescale by std) for the full trajectory are listed after the model names.
Figure 11: Comparing outputs from different models. The target is the first 6 output steps from PDEArena-Incomp dataset (particle density channel). For each model(each row), we display the difference between target and model output. Errors (rescale by std) for the full trajectory are listed after the model names.
@@
We compare the output of different models in fig.˜9, fig.˜10, and fig.˜11. Figures 12 and 13 present additional visualizations of the VICON model outputs compared to ground truth and baseline predictions, highlighting the qualitative advantages of our approach.

### v5

@@ D.2 Analysis of Checkerboard Artifacts in PDEBench-Comp-HighVis
Training-inference normalization mismatch. During training, context pairs are sampled uniformly across all frames. During inference, context pairs are sampled from initial frames only. As shown in Table˜11, the standard deviations at these two different time ranges are vastly different, which results in normalization mismatch and exacerbates prediction errors in the already-challenging problem. Future work could explore using global statistics from multiple trajectories instead of single-trajectory instance normalization during deployment. Other potential remedies include using overlapping patches or moving patch positions during rollout, though these represent engineering improvements orthogonal to our core contributions.
D.3 Cross-Trajectory Context Sampling
We investigate whether VICON can leverage context pairs from different trajectories governed by the same underlying operator, as suggested in the original ICON framework. Using the PDEArena-Incomp dataset (which provides 32 distinct trajectories per operator), we compare VICON trained with intra-trajectory context only versus a mixture of intra- and cross-trajectory context (50/50 split).
Table 12: Cross-trajectory context experiment on PDEArena-Incomp (normalized RMSE, scaled by std)
Training Setup Inference Setup Avg Error
VICON (trained with intra-traj only) intra-traj context 0.563
cross-traj context 0.609
VICON (trained with mixed context) intra-traj context 0.556
cross-traj context 0.568
As shown in Table˜12, VICON is capable of leveraging cross-trajectory information when trained appropriately. The model trained with mixed context achieves competitive performance in both settings, with only a marginal gap (0.556 vs 0.568) between intra- and cross-trajectory inference.
While these results demonstrate VICON’s flexibility, we maintain that intra-trajectory context is more practical for deployment scenarios. Cross-trajectory sampling requires: (1) a pre-existing database of trajectories, and (2) a retrieval mechanism to ensure sampled trajectories share the same (often unknown) operator. In contrast, our intra-trajectory approach is self-contained and better suited for real-time forecasting where only the current trajectory is available.
D.4 More Results and Visualizations
@@
Step 1 [1e-2] PDEArena-Incomp 11.01 11.01 5.97 6.17
@@
Step 5 [1e-2] PDEArena-Incomp 22.68 20.62 11.70 14.81
@@
Step 10 [1e-2] PDEArena-Incomp 35.67 30.53 20.74 27.89
@@
Last step [1e-2] PDEArena-Incomp 76.77 68.03 65.27 93.52
@@
All average [1e-2] PDEArena-Incomp 56.27 48.50 41.20 55.95
@@
Step 1 [1e-2] PDEArena-Incomp 5.63 5.63 3.02 3.12
@@
Step 5 [1e-2] PDEArena-Incomp 10.38 9.46 5.37 6.79
@@
Step 10 [1e-2] PDEArena-Incomp 14.65 12.55 8.55 11.49
@@
Last Step [1e-2] PDEArena-Incomp 16.50 14.56 14.03 19.47
@@
All average [1e-2] PDEArena-Incomp 16.26 14.31 11.86 15.77
@@
Tables 13 and 14 summarize the relative and absolute $L^{2}$ rollout errors across different timesteps for all evaluated models on the three datasets.
@@
Figure˜5 illustrates VICON’s performance with varying timestep strides ($s_{\text{max}}=1,3,5,7$) across all three datasets, demonstrating the model’s ability to adapt to different temporal resolutions without retraining.
@@
Figure˜6 displays the visualization of TKE fields for ground truth and model predictions, highlighting VICON’s superior ability to preserve critical turbulent flow structures compared to baseline approaches.
@@
Table 15: Summary of Rollout Relative $L^{2}$ Error Metrics (single step, scale by std) for 2 Training Strategies. The best results are highlighted in bold. For separate training (a model on each dataset), each run’s batch size is controlled to be the same as joint training. To ensure fair comparison, we maintain the same batch sizes as joint training for each separate training run while adjusting their training steps to be slightly more than one-third of the joint training duration, ensuring comparable total computational costs.
Rollout Relative $L^{2}$ Error [1e-2] Case Joint Separate
Step 1 PDEArena-Incomp 11.10 9.66
@@
Step 5 PDEArena-Incomp 23.00 24.09
@@
Step 10 PDEArena-Incomp 36.18 41.66
@@
Last Step PDEArena-Incomp 77.81 123.0
@@
All average PDEArena-Incomp 56.26 76.35
@@
Figure˜7 and Table˜15 quantify the performance differences between jointly trained and separately trained models, demonstrating the significant advantages of multi-physics training across all datasets.
@@
Table˜16 summarizes the computational resources required by each method, showing VICON’s advantages in terms of training cost, inference speed, and model parameter count compared to both DPOT and MPP baselines.
@@
Figure˜8 show the rollout results with imperfect temporal measurements.
Figure 9: Comparing outputs from different models. The target is the first 6 output steps from PDEBench-Comp-HighVis dataset (x-velocity channel). For each model (each row), we display the difference between target and model output. Errors (rescaled by std) for the full trajectory are listed after the model names.
Figure 10: Comparing outputs from different models. The target is the first 6 output steps from PDEBench-Comp-LowVis dataset (pressure channel). For each model (each row), we display the difference between target and model output. Errors (rescaled by std) for the full trajectory are listed after the model names.
Figure 11: Comparing outputs from different models. The target is the first 6 output steps from PDEArena-Incomp dataset (particle density channel). For each model (each row), we display the difference between target and model output. Errors (rescaled by std) for the full trajectory are listed after the model names.
@@
We compare the output of different models in Figure˜9, Figure˜10, and Figure˜11. Figures 12 and 13 present additional visualizations of the VICON model outputs compared to ground truth and baseline predictions, highlighting the qualitative advantages of our approach.